OpenStudy (mrhoola):

Suppose the time it takes a data collection operator to fill out an electronic form for a database is uniformly distributed between 1.1 and 2.0 minutes. (d) What is the probability it will take more than 90 seconds to fill out the form? (Round your answer to 3 decimal places.)

OpenStudy (mrhoola):

Is the integral's lower limit 9 and upper limit infinity ?

OpenStudy (mrhoola):

or do you solve for the time it takes to fill out the form : integral lower limit 1.1 and upper limit is 9 , then execute 1- result ??

OpenStudy (anonymous):

change minutes to seconds

OpenStudy (anonymous):

1.1 min = 66 sec 2 min = 120 sec

OpenStudy (anonymous):

f(x) = 1 / ( 120 -66) = 1/54

OpenStudy (anonymous):

$$ \Large \int_{90}^{120} \frac{1}{54} ~dx $$

OpenStudy (mrhoola):

Thank you

OpenStudy (mrhoola):

ok so , the probability it will take less than two minutes to fill out the form? from what you have shown me it should be : \[\int\limits_{1.1}^{2.0} 1/(2.0-1.1) dx\]

OpenStudy (mrhoola):

the solution is 1 and that makes sense because the interval limit is 2 t

OpenStudy (anonymous):

right